# Identifying and Addressing Delusions for Target-Directed Decision Making

### v3

@@ Abstract
We are interested in target-directed agents, which produce targets during decision-time planning, to guide their behaviors and achieve better generalization during evaluation. Improper training of these agents can result in delusions: the agent may come to hold false beliefs about the targets, which cannot be properly rejected, leading to unwanted behaviors and damaging out-of-distribution generalization. We identify different types of delusions by using intuitive examples in carefully controlled environments, and investigate their causes. We demonstrate how delusions can be addressed for agents trained by hindsight relabeling, a mainstream approach in for training target-directed RL agents. We validate empirically the effectiveness of the proposed solutions in correcting delusional behaviors and improving out-of-distribution generalization.
@@ 1 Introduction
One important use of Reinforcement Learning (RL) is to train agents with skills generalizable to target tasks for application, after learning from limited training tasks, where the evaluation and training task distributions may differ (Ada et al., 2024). Despite numerous efforts to bridge the “generalization gap” over discrepancies between the two sets of tasks, generalization abilities of RL agents remain largely unsatisfactory. This means that learned skills become less useful to tasks in practice, greatly limiting the scope of applications and research interests. Recent works attributed this to existing agents’ lack of reasoning abilities to face Out-Of-Distribution (OOD) changes (Di Langosco et al., 2022), as most agents either operate solely based on intuition (including model-free methods and offline model-based methods) (Kahneman, 2017). These suggestions led to the development of agents that reason adaptively in novel situations.
@@
Thus, some decision-time planning agents are developed, which make use of their planning outcomes, which we call targets, to adapt in OOD scenarios (Alver & Precup, 2022). Specifically, targets are self-generated (sub-)goals produced during decision-time planning. The “target-directed” agents make use of generators to propose targets, as well as optional estimators that evaluate the favorability of the targets, to assist solving the given tasks. These processes however provide no inherent assurance to the selected targets.
@@
Delusions are obviously wrong beliefs that some agents hold despite contradictory evidence. Formally, they reflect the inability to reject false beliefs (Kiran & Chaudhury, 2009), which in the RL context translates to false beliefs that are natural results of an agent’s improper learning process. Particularly problematic for target-directed agents, delusional behaviors appear in the form of chasing nonexistent or unreachable targets, or those beyond safety constraints (Bengio et al., 2024), which in-turn limit the agents’ situational understanding, and lead to worse generalization in the target tasks or even safety catastrophes. Delusions stem from the incoordination of the generator and the estimator, respectively: problematic targets should not be generated, and even if they were, they should not be favored (Corlett, 2019). While delusions can have various sources, this paper focuses on the misevaluation by the estimator caused by improper training processes, a combination of inadequate update rules and inappropriate training data distributions.
@@
In this paper, we demonstrate how to preemptively address failure modes induced by delusions in the target tasks by training the estimators properly. With clear diagnoses in the designed environments, we identify types of delusions, and investigate more proper designs of the training processes. We then applied our idea to hindsight relabeling and validated in experiments that our strategies led to better understanding of the patterns of delusions and significant reduction in delusional behaviors, resulting in significant improvements in OOD generalization performance.
@@ 2 Preliminaries
RL & Problem Setting.: An RL agent takes actions to interact with an environment to gain cumulative reward. The interaction may be modeled as a Markov Decision Process (MDP) $\mathcal{M}\equiv\langle\mathcal{S},\mathcal{A},P,R,d,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ are the set of states and actions, $P:\mathcal{S}\times\mathcal{A}\to\text{Dist}(\mathcal{S})$ is the state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $d:\mathcal{S}\to\text{Dist}(\mathcal{S})$ is the initial state distribution, and $\gamma\in(0,1]$ is the discount factor. An agent needs to learn a policy $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ that maximizes the value, i.e. the expected discounted cumulative reward $\mathbb{E}_{\pi,P}[\sum_{t=0}^{T_{\perp}}\gamma^{t}R(S_{t},A_{t},S_{t+1})|S_{0}\sim d]$, where $T_{\perp}$ denotes the time step when the episode terminates. Often, environments are partially observable, which means, instead of a state, the agent receives an observation $x_{t+1}$, with which the agent needs to infer the state from the history.
@@
Figure 1: Target-Directed Frameworks: A generator proposes candidate target(s). An estimator can be used to evaluate and select the favorable among candidates.
@@
Target-Directed RL offers a perspective to better identify delusions. It abstracts existing methods that belong to the intersection of decision-time planning and generalized goal-conditioned RL (Nasiriany et al., 2019; Zadem et al., 2024). The framework emphasizes an algorithmic design, where a generator proposes candidate targets, and an (optional) estimator evaluates them and select one target for (optionally target-conditioned) policies to execute (Dayan & Hinton, 1992; Parr & Russell, 1997). The two components correspond to the belief formation and belief evaluation systems, respectively, whose incoordination causes delusions in the human brain (Kiran & Chaudhury, 2009), as shown in Fig. 1.
@@
Hindsight Relabeling is a mainstay approach to train conditioned policies, estimators, as well as temporally-extended generators. Training target-directed agents with only contemporary targets (the ones being followed) can lead to poor performance (Dai et al., 2021), especially for agents that propose their own targets, where contemporary targets may be low-quality, hard to achieve, or lack in diversity (Moro et al., 2022; Davchev et al., 2021). Hindsight Experience Replay (HER) addresses this by creating a buffer of training data for the learner to sample from and filling it with relabeled transitions, which are augmented with alternative targets (Andrychowicz et al., 2017). HER augments a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ with an additional observation $x^{\odot}$ (or its some encoding), the relabeled target. Each (re)labeled transition constitutes a source-target pair, with one decision point for the current step and another for the relabeled target. Relabeling strategies, which correspond to the sources of $x^{\odot}$, are critical for the performance of HER-trained agents (Shams & Fevens, 2022). Most popular choices are trajectory-level, meaning $x^{\odot}$ comes from the same trajectory as $x_{t}$ and are thus conducted at the end of each episode. These include “future”, where $x^{\odot}=x_{t^{\prime}}$ with $t^{\prime}>t$, and “episode”, with $0\leq t^{\prime}\leq T_{\perp}$.
@@
SSM In most works, failure modes of target-directed agents are often overlooked, possibly due to the complexity of benchmark environments. To better identify the causes, and provide intuitive examples, we craft a set of fully-observable environments based on the MiniGrid-BabyAI framework (Chevalier-Boisvert et al., 2018b; Hui et al., 2020), named SwordShieldMonster (SSM for short). In SSM, the agent moves one step at a time in four directions to navigate fields with episode-terminating, randomly placed lava traps. The density of traps is controlled by a difficulty parameter $\delta$, which guarantees a viable path to success, i.e. acquiring a sparse terminal reward. Agents must collect a sword and a shield, both randomly placed, before reaching the “monster” guarding the treasure. Reaching the monster prematurely ends the episode without reward. Visualizations of SSM are in Fig. 2, along with examples of delusional behaviors discussed later. The sword and shield are acquired by moving to their respective grid cells, and agents cannot revert the process. This introduces temporary unreachability in the state structure, as non-terminal states are not fully traversable from one another. Semantically, this segments SSM states into four equivalence classes, formed by two binary statuses: the agent’s possession of the sword and shield. For example, $\langle 0,1\rangle$ denotes sword not acquired, shield acquired.”
@@ 3 Delusions in Target-Directed RL
Delusions can lead to agent chasing outcomes that damage the agents’ generalization in the target tasks or those with catastrophic consequences. We would like to identify delusions, develop corresponding mitigating strategies, s.t. through learning on the training tasks, the agents can identify and avoid delusional behaviors in the target tasks. We look into the two high-level components of target-directed frameworks, i.e. the generator and the estimator, for the sources of delusions and the possible causes.
@@
Figure 2: Visualizations of Delusional Behaviors in SSM: The agent is marked as a red triangle, lava traps as orange squares, and monsters as green loops. In both cases, the estimators, lacking exposure to problematic targets during training, failed to estimate their reachability. This led to the delusion that these problematic targets (yellow dots), seemingly close to every other state, can be used as proxies to reach the distant task goal quickly.
3.1 Generator Delusions
Most learned generators are subjected to unwanted generalization, and hence will inevitably generate problematic targets, a behavior sometimes referred to as hallucination (Jafferjee et al., 2020). While hallucination can be reduced, they are generally unavoidable. In nature, hallucinations are symptoms of delusions, as the issue of unwanted generalization is unlikely to be resolved by the agent autonomously. While delusional behaviors are not all from problematic targets, generator hallucinations are a major source of risks related to delusions.
Generators are often implemented as neural networks and trained in various ways, where direct causes of problematic target generation could be complicated and hard to enumerate. Hence, we focus on identifying the types of problematic targets based on their characteristics.
@@ 3.1 Generator Delusions
The first type of problematic targets includes those that do not correspond to valid states in the task MDP. They can be classified into: 1) Invalid: semantically invalid targets for the task, such as a target singleton of an imagined SSM observation without the agent position; 2) Impossible: semantically valid but unreachable targets like a SSM target cell surrounded by lava traps, (such as in Fig. 2 a). G.3.1 targets only cause delusional behavior if the estimator produces matching Type E.3.2.1 delusions, to be discussed later. Imperfectly generated targets corresponding to real states may not be problematic, as estimators may filter out imperfections, possibly using a state encoder.
@@
The second type includes those that correspond to states in the task MDP, but cannot be fulfilled from the current state. Unlike G.3.1, these targets are only temporarily inappropriate. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$. Naïvely applying relabeling strategies, e.g.episode”, to generator training, can be a major cause of G.3.1.1 hallucinations. Similarly, G.3.1.1 may not result in delusional behaviors if the matching Type E.3.2.2 delusions are not produced by the estimator.
G.3.1.1 targets may appear more frequently if we do not intentionally train the generators to avoid generating them. For example, a conditional target generator which learns from “episode” will be more likely to produce G.3.1.1 targets, such as the one visualized in Fig. 2 b). This intuition applies to training methods beyond HER as well. Importantly, note that G.3.1.1 targets do not exist in all MDPs.
@@
3.2 Estimator Delusions
An estimator evaluates the favorability of proposed targets, based on some relationship between the current state and the targets, which often involve assessing the implementability of targets, represented as, e.g., distances, etc. Acting as a firewall, the estimator of a target-directed agent can enable the selection among the proposed targets, bearing the potential to filter out hallucinated targets. Agents without estimators, e.g. Hafner et al. (2022), must accept proposed targets unconditionally, and thus are at significant risk of chasing problematic targets.
Incorporating an estimator does not mean delusions addressed, as its delusional beliefs about targets can directly cause delusional behaviors. For an estimator to learn to reject delusions, we need 1) effective learning rules that can punish the favorability of problematic targets, and 2) training data that can counteract delusions through the learning rules. A lack of either can result in blind spots of knowledge about targets, delusional or non-delusional.
Update Rules: we expect estimators to learn to reject problematic, undesirable or unsafe targets, by evaluating them with low favorability. However, such expectation may not be achieved, if the estimators’ update rules do not have the key characteristic that punishes unachievable / unsafe targets, when these targets are sampled during training for the estimator to learn from.
Training Data Distribution Even with capable update rules, delusions can still arise from the lack of exposure to data counteracting the estimators’ false beliefs. As shown in Fig. 1, training data are created by organizing mechanisms based on experienced trajectories, e.g. hindsight relabeling for HER-trained agents. Despite these mechanisms can lead to more diversity in training data, they are often not properly set to counter delusions. Take hindsight relabeling as an example: 1) Certain relabeling strategies intrinsically cause delusions. For instance, “future” only relabels future observations, thus only exposes a learner to reachable targets, confusing the estimator when a “past” target is evaluated at decision-time; 2) Trajectory-level relabeling can also be problematic. Short trajectories, common in many training procedures, only cover limited portions of the state space and prevent estimators from learning to evaluate distant targets. Short trajectories can be a product of experimental design (initial state distributions, Maximum Episode Lengths, MELs (Erraqabi et al., 2021)) or innate environmental characteristics (density of terminal states). It is worth noting that delusions caused by improper training data (organizing mechanisms) also extend to cases like few-shot generalization.
With these root causes, we identify types of estimator delusions based on the targets they evaluate. We use the identifiers E.3.2, E.3.2.1, and E.3.2.2 for their resulting delusional behaviors.

### v4

@@ Abstract
Target-directed agents utilize self-generated targets, to guide their behaviors for better generalization. These agents are prone to blindly chasing problematic targets, resulting in worse generalization and safety catastrophes. We show that these behaviors can be results of delusions, stemming from improper designs around training: the agent may naturally come to hold false beliefs about certain targets. We identify different types of delusions via intuitive examples in controlled environments, and investigate their causes and mitigations. With the insights, we demonstrate how we can make agents address delusions preemptively and autonomously. We validate empirically the effectiveness of the proposed strategies in correcting delusional behaviors and improving out-of-distribution generalization.
@@ 1 Introduction
An important use of Reinforcement Learning (RL) is to learn skills generalizable for application, after training from limited training tasks. Despite numerous efforts to bridge the “generalization gap” between performance during training and evaluation, generalization abilities of RL agents remain largely unsatisfactory. Recent works attributed this to existing agents’ lack of reasoning abilities to face Out-Of-Distribution (OOD) changes (Di Langosco et al., 2022).
@@
Thus, a certain type of decision-time planning agents are developed, which make use of their more adaptive planning outcomes, which we call targets, to direct its behaviors to generalize better in novel situations (Alver & Precup, 2022). For instance, these targets can help decompose a complicated given task into smaller and more familiar steps. In this paper, we use targets to denote (sub-)goals produced by the agents themselves during decision-time planning. These “target-directed” agents make use of generators to propose candidate targets, as well as optional estimators to evaluate the favorability of the candidates, to assist the selection of a target to commit to.
@@
While target-directed agents were supposed to be able to generalize well OOD, they are often observed to be blindly chasing unreachable or unsafe targets (Bengio et al., 2024). While existing literature has largely neglected the investigations into these behaviors, we introduce the readers to a new perspective to show that these agents are likely (unintentionally) designed to be delusional. Unlike hallucinations, delusions are obviously wrong beliefs that an agent holds, and reflect an inability to reject false beliefs (Kiran & Chaudhury, 2009). These in the RL context correspond to false beliefs that are natural results of an agent’s improper learning process, which limit the agents’ situational understanding and lead to worse generalization or even safety catastrophes. Delusional behaviors stem from the incoordination of the generator and the estimator: problematic targets should not be generated, and even if they were, they should not be favored (Corlett, 2019).
@@
This paper focuses on necessary designs for an agent to address its delusions autonomously and preemptively: 1) incorporation of an estimator, which enables the rejection of problematic targets; 2) adequate learning rules with which false beliefs about targets can be counteracted and 3) proper training data that include those counteracting delusions. With clear diagnoses in the designed environments, we identify types of delusions, and investigate more proper designs of the training processes. We then applied our idea to hindsight relabeling, a popular and simple form of target-directed RL training and validated in experiments that our strategies led to lower delusion-related errors and significant reduction in delusional behaviors, resulting in significant improvements in OOD generalization performance.
@@ 2 Preliminaries
RL & Problem Setting.: An RL agent takes actions to interact with an environment to gain rewards. The interaction may be modeled as a Markov Decision Process (MDP) $\mathcal{M}\equiv\langle\mathcal{S},\mathcal{A},P,R,d,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ are the sets of states and actions, $P:\mathcal{S}\times\mathcal{A}\to\text{Dist}(\mathcal{S})$ is the state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $d:\mathcal{S}\to\text{Dist}(\mathcal{S})$ is the initial state distribution, and $\gamma\in(0,1]$ is the discount factor. An agent needs to learn a policy $\pi:\mathcal{S}\to\text{Dist}(\mathcal{A})$ that maximizes the value, i.e. the expected discounted cumulative reward $\mathbb{E}_{\pi,P}[\sum_{t=0}^{T_{\perp}}\gamma^{t}R(S_{t},A_{t},S_{t+1})|S_{0}\sim d]$, where $T_{\perp}$ denotes the timestep when the episode terminates. Often, environments are partially observable, which means, instead of a state, the agent receives an observation $x_{t+1}$, with which the agent needs to infer the state from the history.
@@
Figure 1: Target-Directed Framework: A generator proposes candidate target(s). An estimator can be used to evaluate and select the favorable among candidates.
@@
Target-Directed RL offers a perspective to better identify delusions, inspired by delusion-related research in psychiatry (Kiran & Chaudhury, 2009). It abstracts existing methods that belong to the intersection of decision-time planning and generalized goal-conditioned RL (Zadem et al., 2024). The framework emphasizes an algorithmic design, where a generator proposes candidate targets, and an (optional) estimator evaluates them and select one target for policies to execute (Dayan & Hinton, 1992). The two components correspond to the belief formation and belief evaluation systems, respectively, whose incoordination causes delusions in the human brain, as shown in Fig. 1.
@@
Source-Target Pairs & Hindsight Relabeling Essentially, target-directed agents are trained on “source-target pairs”, which are organized training data that could let the agents learn about the relationship between one state and another. Training target-directed agents with only contemporary targets (the ones being followed) can lead to poor performance (Dai et al., 2021), since contemporary targets may be low-quality, hard to achieve, or lack in diversity (Moro et al., 2022; Davchev et al., 2021). Hindsight Experience Replay (HER) is proposed to be an experience organization mechanism transforming collected experiences into relabeled transitions augmented with alternative targets (Andrychowicz et al., 2017). HER augments a transition $\langle x_{t},a_{t},r_{t+1},x_{t+1}\rangle$ with an additional observation $x^{\odot}$ (or its encoding), the relabeled target, creating a source-target pair, with one decision point for the current step and another for the relabeled target. Relabeling strategies, which correspond to how $x^{\odot}$ is selected, are critical for the performance of HER-trained agents (Shams & Fevens, 2022). Most popular choices are trajectory-level, meaning $x^{\odot}$ comes from the same trajectory as $x_{t}$. These include “future”, where $x^{\odot}=x_{t^{\prime}}$ with $t^{\prime}>t$, and “episode”, with $0\leq t^{\prime}\leq T_{\perp}$.
@@
SSM In existing works, failure modes of target-directed agents are often overlooked, possibly due to a lack of access to environmental ground truth in benchmark environments. To identify the causes, and provide intuitive examples, we craft a set of fully-observable environments based on the MiniGrid-BabyAI framework (Chevalier-Boisvert et al., 2018b; Hui et al., 2020), named SwordShieldMonster (SSM for short). In SSM, the agent moves one step at a time in $4$ absolute directions to navigate fields with episode-terminating, randomly placed lava traps. The density of traps is specified by a difficulty parameter $\delta$, while guaranteeing a viable path to success, i.e. acquiring a sparse terminal reward. Agents must collect a sword and a shield, both also randomly placed, before reaching the “monster” guarding the treasure. Reaching the monster prematurely ends the episode without reward. Visualizations of SSM are in Fig. 2, with examples of delusional behaviors discussed later. The sword and shield are picked up by moving to their respective grid cells, and agents cannot drop them. Such design introduces temporary unreachability in the state structure, as non-terminal states are not fully traversable from one to another. Semantically, this segments SSM states into $4$ equivalence classes, formed by $2$ binary indicators: the agent’s possession of the sword and shield. For example, $\langle 0,1\rangle$ denotes sword not acquired, shield acquired”. Despite SSM’s simplicity in terms of observation space, a similar set of environments much simpler than SSM was already shown to be a significant challenge for state-of-the-art methods due to the OOD challenges (Zhao et al., 2024).
@@ 3 Delusions in Target-Directed RL
To ultimately make the agents able to address delusions autonomously, we first need to identify where the problems reside. We first look into the generator to find its contributions to the delusional behaviors, and then locate delusions and their causes in the estimator.
@@
Figure 2: Visualizations of Delusional Behaviors in SSM: The agent location is marked as a red triangle, lava traps’ as orange squares, and monsters’ as green loops. In both a) & b), the estimators, lacking exposure to problematic targets during training, failed to estimate their reachability. This led to the delusion that these problematic targets (whose location are marked with yellow dots), seemingly close to every state, can be used as proxies to reach the distant task goal almost instantly.
3.1 Generator’s Role in Delusional Behaviors: Bad Targets
Most learned generators are subjected to unwanted generalization, thus will inevitably generate problematic targets, a behavior sometimes characterized as hallucination (Jafferjee et al., 2020). While hallucination can be reduced to a degree, they are generally unavoidable, especially in OOD scenarios. Despite that not all delusions are caused by problematic targets, generator hallucinations pose a major risk to safety.
Generators are implemented in various ways, thus it is ineffective to enumerate the detailed causes of problematic target generation. We skip these discussions and directly focus on a categorization of problematic targets, based on an often neglected perspective - the temporal relationship inside the source-target pairs. This leads to the following two disjoint types of problematic targets:
@@ 3.1 Generator’s Role in Delusional Behaviors: Bad Targets
The first type of problematic targets includes those that do not correspond to valid states in the task MDP. They can be classified into: 1) Invalid: semantically invalid targets for the task, such as a target SSM observation generated without the agent’s location; 2) Impossible: semantically valid but unreachable targets, e.g. an SSM target cell surrounded by lava traps, (such as in Fig. 2 a). G.3.1 targets only cause delusional behavior if the estimator produces matching E.3.2.1 delusions, to be discussed later. Imperfectly generated targets corresponding to valid states may not be problematic, as estimators may filter out imperfections, possibly using a state encoder.
@@
The second type includes targets that correspond to valid states in the task MDP, but cannot be fulfilled from the current state. Unlike G.3.1, these targets are only temporarily inappropriate and could be perfectly good targets if the agent were currently to be in a different state. There are some notable subtypes, such as: 1) Irreversible: targets once reachable but are now blocked due to a past transition. For instance, after acquiring the sword in SSM, the agent transitions from class $\langle 0,0\rangle$ to $\langle 1,0\rangle$, sealing off access to $\langle 0,0\rangle$ or $\langle 0,1\rangle$; 2) Segregated: targets that can only be reached from specific initial states that are not available to the agent. For example, an agent spawned in class $\langle 1,0\rangle$ cannot reach states in $\langle 0,0\rangle$ or $\langle 0,1\rangle$.
G.3.1.1 targets are often overlooked in literature, since hallucinations are mostly discussed in contexts without temporal progression, and also they do not exist in all MDPs. G.3.1.1 can appear more frequently if we do not intentionally train the generators to avoid generating them. For example, a conditional target generator which learns from “episode” will be more likely to produce G.3.1.1 targets, such as the one in Fig. 2 b). This applies to training beyond HER as well.
@@
3.2 Estimator’s Role in Delusional Behaviors: Bad Evaluation
An estimator evaluates the favorability of proposed targets, based on some relationship between the current states (sources) and the candidates (targets), which often involve assessing the implementability of targets, represented as, e.g., distances, etc. Acting as a firewall, the estimator of a target-directed agent can enable the selection among the candidates, enabling the potential to filter out hallucinated targets. Agents without estimators, e.g. Director (Hafner et al., 2022), must accept proposed targets unconditionally, and thus are at significant risk of chasing problematic targets.
Incorporating an estimator does not mean delusions addressed, as the estimators’ own delusional beliefs about targets can directly cause delusional behaviors. A lack of either of the following two aspects can result in blind spots of knowledge about targets, i.e. delusions regarding both delusional or non-delusional targets.
Update Rules: we expect estimators to learn to reject problematic, undesirable or unsafe targets, by evaluating them with low favorability. However, such expectation may not be achieved, if the estimators’ update rules do not have the characteristic that (continuously) punishes unachieved / unsafe targets, when these targets are sampled during training for the estimator to learn from.
Training Data Even with capable update rules, delusions can still arise from the lack of exposure to data counteracting the estimators’ false beliefs. This can often be caused by improper experience organization mechanisms, which are responsible for transforming the experienced interactions with environments into training data, as shown in Fig. 1. Despite these mechanisms like HER can lead to more diversity in training data, they are often not capable of avoiding delusions: 1) Certain relabeling strategies naturally cause delusions. For instance, “future” only relabels with future observations, thus only exposes a learner to future reachable targets, leaving the estimator to guess when a “past” target is proposed at decision-time; 2) Trajectory-level relabeling can also be problematic. Short trajectories, common in many training procedures, only cover limited portions of the state space and prevent estimators from learning about distant targets, leaving estimators to guess when a distant target is generated at decision-time. Short trajectories can be a product of experimental design (initial state distributions, Maximum Episode Lengths, MELs (Erraqabi et al., 2021)) or innate environmental characteristics (density of terminal states). It is worth noting that delusions caused by improper training data (organizing mechanisms) also extend to cases like few-shot generalization.
With these necessary conditions, we identify types of estimator delusions based on the targets they evaluate. We use the identifiers E.3.2, E.3.2.1, and E.3.2.2 for their resulting delusional behaviors.